6. Heart failures are due to either natural occurrence (87%) or outside factors (13%). Outside factors are related to induced substances (73%) or foreign objects (27%). Mat ural occurrences are caused by arterial blockage (56%), disease (27%), and infection (eg, staph infection) (17%). (a) Determine the probability that a failure is due to an induced substance. (b) Determine the probability that a failure is due to disease or infection.
Answer #1

6. (a) Supposing there are n occurrences of failure. Then, due to outside factors, there would be 0.13n occurrences. Out of those, there would be 0.13n0.73 occurrences due to induced substances. Hence, the probability that a failure is due to induced substance would be 0.13n 0.73 0.13 0.73-0.0949 or 9.49%.

The event of failure due to induced substance would be the event of failure due to induced substance and outside factor. This can also be done as for we have PlOF) = 0.13 (probability of outside factor) and P(ISOF) = 0.73 (probability of induced factor given failure is due to outside factor). Hence, as P(IS cap OF) = P(IS | OF)*P(OF) , we have P(ISnOF) = 0.73 * 0.13= 0.0949 or 9.49%, which would be the probability of failure due to induced substances.

(b) Again, supposing that there are n occurrences of failure. Then, due to natural occurrences, there would be 0.87n events. Among these, there would be 0.87n*0.27 cases due to disease and 0.Sin 0.1 due to infection. The probability that failure is due to disease would be rac{0.87n*0.27}{n} = 0.2349 , and probability that failure is due to infection would be rac{0.87n*0.17}{n} = 0.1479 . The probability that the failure is due to disease or infection would be 0.2349 0.1479 0.3525 or 38.28%.

This can also be done as we have P(NO) = 0.87 , P(D|NO) = 0.27 and P(I|NO) = 0.17 . Hence, we have P(D cap NO) = 0.87 * 0.27 and P(I cap NO) = 0.87 * 0.17 . The probability that failure is due to D or I would be P(D cap NO) + P(I cap NO) = 0.87 * 0.27 + 0.87 * 0.17 = 0.3828 or 38.28%.